OpenStudy (anonymous):

How do I set c+t+v=366, 2t=v &186+v=c as a matrices?

jimthompson5910 (jim_thompson5910):

First, get all equations in standard form 2t=v ---> 0c + 2t - v = 0 186+v=c ---> c + 0t -v = 186 Now collect the coefficients and write them down in matrix form to get 1 1 1 366 0 2 -1 0 1 0 -1 186

OpenStudy (anonymous):

Thanks but when it says 186 more c than v, it means "196+v=c" right?

jimthompson5910 (jim_thompson5910):

I think you mean 186+v = c, if so, you're right but you have to get the equations into standard form

OpenStudy (anonymous):

cant i just do that once I plug in the answer? Cause that up there looks like something my teacher did on the board but I turned in. :/

jimthompson5910 (jim_thompson5910):

but if you want to place those equations in matrix form, you have to get them all in standard form first from there, you can row reduce to get the answer

OpenStudy (anonymous):

Okay, thanks. :) Since you're good with this, can you please help me with the other one.

jimthompson5910 (jim_thompson5910):

sure, what's the question

OpenStudy (anonymous):

Thanks, okay A theater group will perform 4 shows of its latest production. Each ticket for Friday or Saturday evening will cost $10. Each ticket for Saturday and Sunday matinee will cost $7. Write a matrix to organize this information

jimthompson5910 (jim_thompson5910):

Are the totals given as well?

OpenStudy (anonymous):

No.

jimthompson5910 (jim_thompson5910):

Let x = # of tickets sold for Friday evening y = # of tickets sold for Saturday evening z = # of tickets sold for Saturday matinee w = # of tickets sold for Sunday matinee So because "each ticket for Friday or Saturday evening will cost $10", we know that 10x+10y = c where c is the total cost for the evening shows and since "Each ticket for Saturday and Sunday matinee will cost $7", we can say that 7z+7w = d where d is the total cost for the matinee shows Now convert each equation to standard form to get 10x+10y+0z+0w = c 0x+0y+7z+7w = d So the system converts to the following matrix 10 10 0 0 c 0 0 7 7 d

jimthompson5910 (jim_thompson5910):

unfortunately, we cannot solve this because we do not have enough info

OpenStudy (anonymous):

Thank you. :)